Jon is 14 years older than Philip. 2 years from now, Jon will be 3 times as old as Philip. How old is each now?
Answer by lwsshak3(11628)

Jon is 14 years older than Philip. 2 years from now, Jon will be 3 times as old as Philip. How old is each now?

Let x = Philip's present age
x+14 = Jon's present age
two years from now
Philip's age = x+2
Jon's age = (x+14)+2 = x+16
as stated, Jon will be 3 times as old as Philip
3(x+2) = (x+16)
3x+6 = x+16
2x=10
x=5
Philip's present age = 5
Jon's present age = 5+14=19
two years from now,
their ages will be 7 & 21 respectively